A 220 volt, 1000 watt bulb is connected across a 110 volt mains supply. The power consumed will be
(A) 750 watt
(B) 500 watt
(C) 250 watt
(D) 1000 watt

Grade:12

2 Answers

Bhaskar Gupta
6 years ago
P = V2/R  
If V = 220 V we have

1000 W = 2202/R
R = 2202/1000 Ω = 48.4 Ω. This is the resistance of the bulb.

When V = 110 V, power consumed =  V2/R = 1102 /48.4 = 250 W.
 
ankit singh
3 years ago
.here is your answer......
P= V2/R  
If V = 220 V we have
1000 W = 2202/R
R = 2202/1000 Ω = 48.4 Ω. This is the resistance of the bulb.
When V = 110 V, power consumed =  V2/R = 1102 /48.4 = 250 W........